the area of a rectangle wall of a barn is 20 feet. its length is 6 feet longer than twice its width. find the length and width of the wall of the barn
Answer by stanbon(75887)

the area of a rectangle wall of a barn is 20 sq feet.
its length is 6 feet longer than twice its width.
find the length and width of the wall of the barn.
--------------------
width:: x
length:: 2x+6
------
Equation:
length * width = 20 sq ft
---
x(2x+6) = 20
2x^2 + 6x - 20 = 0
x^2 + 3x - 10 = 0
Factor:
(x+5)(x-2) = 0
Positive solution::
x = 2 ft (width)
2x+6 = 2*2+6 = 10 ft (length)
---------------
Cheers,
Stan H.
---------------